Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
bonnand, pillant, mwas, gouaty, letendre, masserot, pacaud. - 21:39 Tuesday 09 July 2019 (46309) Print this report
DET recovery - investigations on faulty camera

Today we open SDB2 at the beginning of the afternoon to adress the issue of the faulty B1 photodiode (B1_PD2).

After putting the bench in-air, we first tried to open the shutter it did not work.

To check if the issue is coming from the PD shutter or the driving of the shutter we swapped the 2 cables of the 2 B1 photodiodes.

It was possible to open the shutter from B1_PD1 with both cable (i.e both channel of the service photodiode mezzanine) and B1_PD2 did not open in any case. Therefore the issue is with the PD shutter and NOT the service photodiode mezzanine.

So it was decided to put the working photodiode (SN07) on the B1_PD2 spot and to swap the 2 beamsplitter in order to have all the power on that photodiode. We put one of the old B1 photodiode (SN09, 16 mA preamplifier) on the B1_PD1 spot.

This photodiode only see a very small fraction of the B1 power (0.50 x 0.001 = 0.0005 of B1). It is mainly there for balancing purpose.

We also replaced B1s2 camera which was not working for several weeks.

After a check of the rough alignment of B1_PD2, by letting a low amount of light getting through the 2 OMCs we closed the tower and started the vacuum.

Finally we had to reconfigure the DAQboxes of SDB2 because the timing was wrong (this happens when blocking and working on the bench). After some time and effort we managed with the remote help of Nicolas Letendre, Alain Masserot and Emmanuel Pacaud to reconfigure the DAQboxes properly.

Michal Was from remote finely aligned the beam on B1_PD2 with the SDB2 bench angular and position control closed.

The recovery of the ITF by Julia Casanueva and Marco Vardaro (from remote also) started around 19:30 LT.

So far Low Noise 1 is reached without troubles.

 

In conclusion, the situation on SDB2 is the following, there is one 1 HQE photodiode (B1_PD2) looking at the B1 beam with an HR mirror in front of it.

Comments to this report:
bonnand - 10:09 Wednesday 10 July 2019 (46313) Print this report

There is a mistake in the title.

It should be "DET recovery - investigations on faulty photodiodes"

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×